OpenStudy (goformit100):

Shabana wants to put a lace border all around a rectangular table cover (Fig 10.3), 3 m long and 2 m wide. Find the length of the lace required by Shabana.

OpenStudy (goformit100):

@Rancd

OpenStudy (anonymous):

10 m long

OpenStudy (kirbykirby):

A rectangle has the formula 2(L+h) for the perimeter (L = length, h = height), so 2(2+3) = 2(5)=10